Mike's insulin sensitivity factor is 40 and his insulin to CHO ratio is 1 to 12. What is the correct pre meal rapid acting insulin dose for Mike to take based on the following? Round to nearest whole # :

Pre meal blood glucose:  67 mg/dl

Lunch : 2 starch exchanges, 5 meat exchanges, 2 fat exchanges, 1 fruit exchange
Answer #1

Starch = 15 gms CHO, So 2 starch contains 30 gms of CHO

Fruit = 15 gms of CHO

Meat and fat exchanges does not contain CHO.

Therefore, CHO for lunch = 15 +30 = 45 gms

Insulin : CHO = 1: 12

So, Pre meal insulin dose = 45 / 12 = 4 units